Calculate the number of grams-formula weights in 50 g of NaCl


1
Expert's answer
2021-03-03T01:32:30-0500

n = m/M

M (NaCl) = 58.44 g/mol

n (NaCl) = 50 / 58.44 = 0.86 mol